OpenStudy (anonymous):

Evaluate: 10 - 64 ÷ 8 + 42

OpenStudy (anonymous):

44

OpenStudy (angela210793):

10-8+42=44

OpenStudy (anonymous):

The 42 Is Actually 4 To The Second Power

OpenStudy (anonymous):

wait no. addition goes before subtraction. 10-8+42=10-50=-40 pemdAs